OpenStudy (anonymous):

Solve -2x + 1 -7 and describe the graph of the solution. Answer x 4; open circle on 4, shading to the left x 4; closed circle on 4, shading to the left x 4; open circle on 4, shading to the right x 4; closed circle on 4, shading to the right

OpenStudy (anonymous):

I think there should be another symbol ????

OpenStudy (anonymous):

Which is it supposed to be? -2x + 1 > -7 -2x + 1 < -7 -2x + 1 >= -7 -2x + 1 <= -7

OpenStudy (anonymous):

greater than or equal to >= Less that or equal to <=

OpenStudy (anonymous):

x<=

OpenStudy (anonymous):

open means greater/less than closed means greater/less or equal to the left means greater to the right means less (for this problem)

OpenStudy (anonymous):

-2x + 1 <= -7 -1 -2x <= -8 /-2 flip sign x >= 4

OpenStudy (anonymous):

ur question is fully wrong one

OpenStudy (anonymous):

Follow @nsutton's advice on how to graph it.

OpenStudy (anonymous):

for solving this.it shoud be a equation or should contain inequality sign

OpenStudy (anonymous):

its an inequality

OpenStudy (anonymous):

x >= 4 closed circle; shading to right Hope I helped